# What is a Social Media Bot? A Global Comparison of Bot and Human Characteristics

### v1

@@ Introduction
In essence, a bot is a computer algorithm. As an algorithm, a bot is neither bad or good. It is the persona it is afforded to that determines the goodness of its use. We develop a generic definition of a bot. The definition is independent of the use of the bot. The determination of the use of the bot warrants separate treatment beyond this paper. Regardless of whether a bot is used for good or ill, the behavioral characteristics of bots remain the same.
@@
  • User: An automated account that carries out a series of mechanics. A key characteristic of bots is its programmability, which give it its artificial and inauthentic characteristic. Automation is an aspect that has been iterated in all the reference definitions. The key here is that a bot is automated. The model underlying the automation does not matter; any model can be applied equally well to humans and bot. A bot could be built to mimic humans, or it could be built to optimize other functions. For example, some studies describe bots in terms of its mimicry of humans [33, 50], but others observe that bots eventually influence the social space such that humans mimic bots [51, 52].
  • Content: for content creation, distribution, and collection and processing. Bots often generate their content in bulk to distribute a certain ideology [24, 53], such as a good portrayal of their affiliated country [54]. Instances where bots perform content distribution is where the spread fake news and disinformation content [28, 55, 56], or when they distribute general news information [57]. Bots in the form of web crawlers and scrapers download and index data from social media in bulk [58], and sometimes process the data to perform functions like analyzing sentiment of opinions [59].
  • Relationships: for relationship formation and dissolution. In other words, forming a relationship online means to connect with other users via post-based (i.e., retweet, mention) or user-based (i.e., following, friend) mechanisms. Dissolving a relationship means to destroy a connection by forcing users to leave a community. Bots are an actively form and destroy relationships on social media platforms. An example of the formation of post-based relationship is the active amplification of narratives. This technique is mostly employed in the political realm where the bots retweet political ideology in an organized fashion [24, 60, 61]. User-based relationships can grow through coordinated fake-follower bots, that are used to boost online popularity [62], or can be dissolved through toxic bots that spread hate and directed religious ideologies [63, 40], causing users to break away from the community [64].
@@
Bridging Bot political commentators that aggregate information [16] cross-cultural social marketing, disseminate information across community differences [54]
@@ Results
3(a) presents the differences between cues used by bots and humans. The detailed numeric differences are in the Supplementary Material. This difference is examined overall, and by event. There are consistent differences in the use of cues by bots and humans. For example, across all events, bots use significantly more abusive terms and expletives, and tweet more than humans. On the other hand, humans use more first person pronouns, positive sentiment, and media (i.e., images, videos). Humans tend to quote and reply to tweets, while bots tend to retweet.
@@
Figure 3: Comparison of psycholinguistic overall cue usage (average cue usage per user) by bots and humans across datasets. Green cells show that humans use a larger number of the cue. Red cells show that bots use a larger number of the cue. * indicates there is a significant difference in the usage of the cue between bots and humans.
When a bot retweets a human, its linguistic profile, by definition, is identical to the human’s. The question though, is whether the bots that are sending original tweets match the linguistic profile of those retweeting, or is the linguistic profile different? For the Black Panther and Captain Marvel events (3(b)), we compared the psycholinguistic profile for all tweets, and the original tweets only (i.e., no retweets). In these two events, bots retweet significantly more than humans. However, the bot-human difference between linguistic cue use of the original tweets vs all tweets are rather similar. Only the average tweet reading difficulty and the number of friends are different: in original tweets, humans have higher values; in all tweets, bots have higher values. Therefore, bots have their unique signature when generating new content, but are guaranteed to match human’s content when retweeting the human’s.
Bots construct tweets with cues that can be easily and heavily automated, while humans construct more personal tweets that require higher cognitive processing to create. Such differences shows how bot accounts still use rather rudimentary techniques: hashtag latching using multiple hashtags [40, 27], connecting several users together with increased number of mentions [54] and flooding the zone with lots of tweets tweets of their desired narratives [24, 84]. More sophisticated communication techniques like having an increased number of media, and more advanced interaction techniques that involve dialogue understanding like increasing the number of replies and quotes, are still left to the humans. In short, bots have not entirely mimicked humans, yet.
@@
Social identity theory depicts how social media users portray their image online, and the community that they want to be associated with [85, 86]. We analyze the difference in the self-presentation of the identities between bots and humans, and the difference between the linguistic cues used by the identities. Across the events, there are consistently a smaller proportion of bots that present with an identity. Overall, 21.4% of bots present an identity, while 27.0% of humans present an identity (see Appendix Table 10). Bots are more likely to obfuscate their identities [87], allowing them to take on different personas to suit their operation requirements [88]. 4(a) presents the top 25 identities by frequency between bots and humans. There is a more exponential drop of the frequency of the use of identities in bot users than in human users, suggesting that bots concentrate their self-presentation on certain identities, mostly the common ones: man, son, fan, lover; while humans have a more varied identity presentation.
We then ask a follow-up question: “How do the same bot/human identities talk about the same topics?" We compare the use of topic frames per identity for the most frequent identity affiliations in 4(b). This plots the percentage difference of the use of framing cues (Family, Gender, Political, Race/Nationality, Religion) between bots and humans. This metric compares the use of cues with the human usage as a baseline. Overall, bots converse more aggressively in all topic frames. In particular, bots converse most around societal fault lines: gender, political, race/nationality. These conversations lie on societal fault lines, which could sow discord and chaos [89], therefore such bots are of interest to monitor and moderate. In fact, bots use more gender-based cues. Other research groups have also identified that a disproportionate number of bots that spread disinformation are females [90, 91], and are thus more likely to use gender frames in their posts. Bots tend to converse largely about political topics, regardless of the identity they affiliate with, indicating that a good proportion of bots are deployed for political purposes, either by political parties or by political pundits [16, 68, 26]. Finally, the difference between the usage of topic frames between bots and humans could be due to their vocabulary used. The words used by humans are more varied and mostly not standard dictionary words, while bots are still being programmed with a limited set of vocabulary, as evidenced by the proportion of words identified by the dictionaries in the NetMapper program used. In a similar aspect, chat bot interactions have a more limited vocabulary than human interactions [92].
4(c) presents the average use of topic frames by identity categories. Humans affiliate themselves equally with all identity categories, while bots generally affiliate themselves with racial and political identities. Both bots and humans converse a lot on gender and political issues.
@@
Figure 4: Comparison of identity-related behaviors in bots and humans
@@
Social interactions between users are an indication of the information dissemination patterns and the communication strategies of the users. We calculate the network metrics (total degree, in degree, out degree, density) of the all-communication ego-networks of the users. In the network graphs, the users are nodes, and the links between users represent all communications between the two users (i.e., replies, quotes, mentions, retweets). Table 5 compares the two metrics for bots and humans. Bot ego networks have higher density than ego networks (8.33% more dense), which reflects that the bots have tighter communication structures and form more direct interactions than humans. On average, a bot has 9.66% bot alters and 90.34% human alters, whereas on average a human has 7.31% bot alters and 92.69% human alters. Although bots interact with a higher proportion of bot alters than humans do (32% more bot alters), our findings show that both bots and humans interact more with humans rather than bots in their ego network. By the principle of homophily, it is natural for humans to interact with other humans [93]. However, bots violate the principle of homophily, and instead of interacting with more bots, they interact with more humans. Therefore, bots are actively forming communication interactions with humans, perhaps attempting to influence humans [94].
@@
A common way for bots to be used in political discourse (e.g. elections) is to amplify other users. As an amplifier, bots are the pendants of the user they are amplifying. Therefore, bots appear in star networks in many of the peripheral nodes. A star structure is a network that have a strongly connected core and peripheral networks. This structure is most prominent in bots in political discourse, where core bots create information, and peripheral bots amplify the information through the retweeting mechanic [24]. Humans, on the other hand, are more likely to be part of a tree structure, where one can make out the tiered first- and second-hop interactions. In the same discourse, humans are more likely to be performing many actions, sometimes retweeting other users, sometimes tagging other users and so forth.
@@ Discussion
Through our large scale empirical study, we show that bots and humans have interesting and consistent differences between them. These differences span from their volume, to the linguistic features of their text, to the identities they affiliate with, to their social network structure. These features can be used to characterize a social media bot, and how it differs from humans. We study a huge amount of data dated from 2018 to 2021. These data show consistent differences over the years, which means that while bot technology do evolve, it does not evolve drastically. Moreover, the consistent differences show that there are scenarios where bots can be better than humans, and scenarios where humans can be better than bots. These differences provide insights to how both can be utilized to afford conversations on social media: bots can be used for methodological postings with deliberate selection of hashtags, tweets per hour, and a structured star communication network. Humans can be used for more complex cognitive tasks such as adding media to a post or replying to a post, and for conversing on a larger range of topics [3].
@@
An Artificial Intelligent (AI) system is a machine, or computer system, that can perceive its environment and use intelligence to perform actions to achieve defined goals [96]. The social media bot perceives the digital environment to decide their targets (i.e., users to retweet, users to mention), and intelligently carry out content and interaction mechanics to achieve their goals (i.e., spreading information [56, 28], sowing discord [97, 98], assisting the community [45, 46, 47]). The software-programmed social media bot is an AI algorithm, and thus has potential to be harnessed for social good.
Table 6 lists some recommendations of our results on how bots can be leveraged on for social good, and how they can be regulated. First, given that bots use more retweets and mentions than humans, and have high tweets per hour, bots can be used for menial tasks like announcements and distribution of information. Second, since bots have a star interaction network, they can be used for big announcements like disaster and crisis management without message distortion. A star network sends messages directly through interactions, hence the original message is preserved. However, the human’s hierarchal interaction network will distort the message as it passes through the tiers. Third, bots typically post content that matches their identity, they can be used to provide educational material about topics that people associate with certain profession. For example, a weather news bot can provide weather information. Lastly, since bots use more abusive and expletive terms than humans, instead of regulating toxic language itself, regulation can be focused on disallowing bots to use such toxic language, which would therefore reduce the amount of hyperbole and offense online.
@@
Bot evolution and bot detection are thus a “never-ending clash" [101], and sometimes bot accounts evolve faster than current known bot detection algorithms [68], presenting several opportunities in continual improvement of bot detection algorithms, specifically to be adaptable, faster, and more efficient. The increasing trends of using Large Language Models and Large Vision Models to create generated texts and deepfakes lend bots a helping hand in the construction of more believable narratives. These same generative technology are also used to construct offensive bots for humor [102]. However, current trends reflect that the use of such technologies are not very prevalent, for example, [73] only found one set of such botnet in their study, reflecting that bots are still relying on traditional techniques, likely because such heuristic-based techniques are easier and faster to deploy en masse.
@@
After identifying which users are likely to be bots, one must differentiate the goodness of the bot and its function. This evaluation can be inferred from the bot’s content postings and relationship interactions. However, bots do not fall squarely in a spectrum of goodness; the lines of good and bad bots are blurred. In fact, bots can move between neutral in which they post messages that are not harmful, to bad, where they post politically charged and extremist messages [40, 103]. Herein lies an opportunity to construct a rubric to determine the goodness of the bot; this, though, is a complex task, for there are ethical and societal issues to consider. Bots can change their goodness, too. They may be supporting a certain cause initially, then making a swing to a different stance soon enough. This swing of support was witnessed during the coronavirus pandemic era, and especially so when the bots require little conviction to change allegiances [8]. Another challenge involves identifying the type of bot, which can provide insight towards possible impact of the bot. For example, an Amplifier Bot that intensifies political messages could be intended to sow discord [24, 61].

### v2

@@ Introduction
In essence, a bot is a computer algorithm. As an algorithm, a bot is neither bad or good. It is the persona , or public facade, that the bot is afforded to that determines the goodness of its use. We develop a generic definition of a bot. The definition is independent of the use of the bot. The determination of the use of the bot warrants separate treatment beyond this paper. Regardless of whether a bot is used for good or ill, the behavioral characteristics of bots remain the same.
@@
  • User: An automated account that carries out a series of mechanics. A key characteristic of bots is its programmability, which give it its artificial and inauthentic characteristic. Automation is an aspect that has been iterated in all the reference definitions. The key here is that a bot is automated. The model underlying the automation does not matter; any model can be applied equally well to humans and bot. A bot could be built to mimic humans, or it could be built to optimize other functions. For example, some studies describe bots in terms of its mimicry of humans [33, 50], but others observe that bots eventually influence the social space such that humans mimic bots [51, 52]. The automated nature of the bot gives rise to its systematic nature, which bot detection algorithms leverage through machine learning algorithms. In contrast, human users create accounts for a wide number of reasons, from personal to business to pet accounts, and have very varied and random behavior mechanics.
  • Content: for content creation, distribution, and collection and processing. Bots often generate their content in bulk to distribute a certain ideology [24, 53], such as a good portrayal of their affiliated country [54]. Instances where bots perform content distribution is where the spread fake news and disinformation content [28, 55, 56], or when they distribute general news information [57]. Bots in the form of web crawlers and scrapers download and index data from social media in bulk [58], and sometimes process the data to perform functions like analyzing sentiment of opinions [59]. Humans create and distribute content to create their online personalities, to update their friend groups of their lives, or to promote their businesses [60, 61]. However, humans seldom generate or collect content in bulk; if they do so, they are likely to use a bot to aid them, rendering their account a cyborg [3].
  • Relationships: for relationship formation and dissolution. In other words, forming a relationship online means to connect with other users via post-based (i.e., retweet, mention) or user-based (i.e., following, friend) mechanisms. Dissolving a relationship means to destroy a connection by forcing users to leave a community. Bots are an actively form and destroy relationships on social media platforms. An example of the formation of post-based relationship is the active amplification of narratives. This technique is mostly employed in the political realm where the bots retweet political ideology in an organized fashion [24, 62, 63]. User-based relationships can grow through coordinated fake-follower bots, that are used to boost online popularity [64], or can be dissolved through toxic bots that spread hate and directed religious ideologies [65, 40], causing users to break away from the community [66]. In general, bots form and dissolve automated relationships towards some single-purpose goal, while humans form and dissolve relationships organically, sometimes based on real-life encounters.
@@
Bridging Bot political commentators that aggregate information [16] disseminate information to instigate anger across social, cultural, community differences [54]
@@ Results
We used a suite of multidisciplinary methods for our analysis. This is illustrated in Figure 2. We integrated machine learning classification, social network analysis and linguistic feature extraction in a robust hybrid methodology that bridged computational, linguistic and network sciences. We analyzed the behavior of bots and humans along four axes. We analyzed the behavior of social media users along four axes. The first is classifying users into bot or human using a machine learning model called BotHunter. The second was linguistic feature analysis of tweets and user metadata. These information were extracted from the NetMapper software, then averaged across user type (i.e., bot vs human) and event type, and compared using statistical tests. The third is the self-presentation of social identities in the user metadata. These identities are derived from matching with an occupation census. In this third analysis, we also extracted topic frames that were written in the tweet texts using the NetMapper software. Thereafter, we correlated the social identities with the topic frames, linking together user presentation with user writing. Fourth, we construct interaction networks of users using the software ORA, and analyzed the influence of bots and humans in terms of their position and structure of their ego networks. Further details about our methodology and the implementation can be found in the Supplementary Material.
Figure 2: Illustration of multidisiplinary methods used for our analysis
@@
4(a) presents the differences between cues used by bots and humans. The detailed numeric differences are in the Supplementary Material. This difference is examined overall, and by event. There are consistent differences in the use of cues by bots and humans. For example, across all events, bots use significantly more abusive terms and expletives, and tweet more than humans. On the other hand, humans use more first person pronouns, positive sentiment, and media (i.e., images, videos). Humans tend to quote and reply to tweets, while bots tend to retweet.
@@
Figure 4: Comparison of psycholinguistic overall cue usage (average cue usage per user) by bots and humans across datasets. Green cells show that humans use a larger number of the cue. Red cells show that bots use a larger number of the cue. * within the cells indicates there is a significant difference in the usage of the cue between bots and humans at the $p<0.05$ level.
When a bot solely retweets a human, its linguistic profile, by definition, is identical to the human’s. The question though, is whether the bots that are sending original tweets match the linguistic profile of those retweeting, or is the linguistic profile different? For the Black Panther and Captain Marvel events (4(b)), we compared the psycholinguistic profile for all tweets, and the original tweets only (i.e., no retweets). In these two events, bots retweet significantly more than humans. In general, the bot-human difference between linguistic cue use of the original tweets vs all tweets are rather similar. However, the average tweet reading difficulty and the number of friends are different: in original tweets, humans have higher values; in all tweets, bots have higher values. Therefore, bots have their unique signature when generating new content (i.e., in original tweets), but are guaranteed to match human’s content when retweeting the human’s.
Bots construct tweets with cues that can be easily and heavily automated, while humans construct more personal tweets that require higher cognitive processing to create. This shows in 4(a), where bots use more semantic cues, while humans use more emotion cues. For metadata cues, bots have more tweets and tweets per hour, while humans engage more with the conversation through replies and quotes. Such differences show how bot accounts still use rather rudimentary techniques: hashtag latching using multiple hashtags [40, 27], connecting several users together with increased number of mentions [54] and flooding the zone with lots of tweets tweets of their desired narratives [24, 86]. More sophisticated communication techniques like having an increased number of media, and more advanced interaction techniques that involve dialogue understanding like increasing the number of replies and quotes, are still left to the humans. In short, bots have not entirely mimicked humans, yet.
@@
Social identity theory depicts how social media users portray their image online, and the community that they want to be associated with [87, 61]. We analyze the difference in the self-presentation of the identities between bots and humans, and the difference between the linguistic cues used by the identities. Across the events, there are consistently a smaller proportion of bots that present with an identity. Overall, 21.4% of bots present an identity, while 27.0% of humans present an identity (see Supplementary Material Table 10). Bots are more likely to obfuscate their identities [88], allowing them to take on different personas to suit their operation requirements [89]. 5(a) presents the top 25 identities by frequency between bots and humans. Overall, bots affiliated with a more limited set of identities (n=869 vs n=908 unique identities). The distribution of the identities for each dataset is presented in the Supplementary Material Tables 3 and 4. There is a larger drop in the frequency of the use of identities in bot users than in human users. For example, between the 3rd and 4th most frequently used identities, bots show a 33.2% decrease in the use of the 4th most frequent identity compared to the 3rd most frequent identities in bots, compared to a 30.0% decrease for humans. This observation suggests that bots concentrate their self-presentation on certain identities, mostly the common ones: man, son, fan, lover; while humans have a more varied identity presentation.
[…]
5(c) presents the average use of topic frames by identity categories. Humans affiliate themselves equally with all identity categories, while bots generally affiliate themselves with racial and political identities. Both bots and humans converse a lot on gender and political issues.
@@
Figure 5: Comparison of identity-related behaviors in bots and humans. The Methods section explains the derivation of the identity categories and topic frames.
@@
Social interactions between users are an indication of the information dissemination patterns and the communication strategies of the users. We calculate the network metrics (total degree, in degree, out degree, density) of the all-communication ego-networks of the users. In the network graphs, the users are nodes, and the links between users represent all communications between the two users (i.e., replies, quotes, mentions, retweets). Table 5 compares the two metrics for bots and humans. Bot ego networks have higher density than human ego networks (8.33% more dense), which reflects that the bots have tighter communication structures and form more direct interactions than humans. On average, a bot has 9.66% bot alters and 90.34% human alters, whereas on average a human has 7.31% bot alters and 92.69% human alters. Although bots interact with a higher proportion of bot alters than humans do (32% more bot alters), our findings show that both bots and humans interact more with humans rather than bots in their ego network. By the principle of homophily, it is natural for humans to interact with other humans [94]. However, bots violate the principle of homophily, and instead of interacting with more bots, they interact with more humans. Therefore, bots are actively forming communication interactions with humans, perhaps attempting to influence humans [95].
@@
A common way for bots to be used in political discourse (e.g. elections) is to amplify other users. As an amplifier, bots are the pendants of the user they are amplifying. Bots echo the ideas, narratives and emotions of these users, enhancing the visibility of the users and their narratives within the social network. Therefore, bots appear in star networks in many of the peripheral nodes. A star structure is a network that have a strongly connected core and peripheral networks. This structure is most prominent in bots in political discourse, where core bots create information, and peripheral bots amplify the information through the retweeting mechanic [24]. Humans, on the other hand, are more likely to be part of a tree structure, where one can make out the tiered first- and second-hop interactions. In the same discourse, humans are more likely to be performing many actions, sometimes retweeting other users, sometimes tagging other users and so forth.
@@ Discussion
Through our large scale empirical study, we show that bots and humans have interesting and consistent differences between them. These differences span from their volume, to the linguistic features of their text, to the identities they affiliate with, to their social network structure. These features can be used to characterize a social media bot, and how it differs from humans.
Our work introduces a valuable dataset containing billions of tweets over four years. This dataset is a valuable resource for studying patterns in social media bot technology. With the current restrictions by social media platforms for data collection, this dataset will be prohibitively expensive and technically not feasible to replicate. X API, this dataset will be extremely costly to collect. Using the free tier API, this dataset would take 50 million months or 136,986 years to complete collection. The Pro tier API costs allows retrieval of 1 million tweets per month at a cost of \$5000, leading to the total cost of \$25 million over 5000 months, or 416 years for the collection to be completed. Table 11 in the Supplementary Material details the calculations of the number of months and the cost required to replicate this dataset under the 2025 pricing structure.
We study a huge amount of data dated from 2018 to 2021. These data show consistent differences over the years, which means that while bot technology do evolve, it does not evolve drastically. Moreover, the consistent behavioral differences show that there are scenarios where the automated nature of bots is more advantageous, and scenarios where the manual human control and thought is superior. These differences provide insights to how both can be utilized to afford conversations on social media: bots can be used for methodological postings with deliberate selection of hashtags, tweets per hour, and a structured star communication network. Humans can be used for more complex cognitive tasks such as adding media to a post or replying to a post, and for conversing on a larger range of topics [3].
@@
An Artificially Intelligent (AI) system is a machine, or computer system, that can perceive its environment and use intelligence to perform actions to achieve defined goals [97]. The social media bot perceives the digital environment to decide their targets (i.e., users to retweet, users to mention), and intelligently carry out content and interaction mechanics to achieve their goals (i.e., spreading information [56, 28], sowing discord [98, 99], assisting the community [45, 46, 47]). The software-programmed social media bot is an AI algorithm, and thus has potential to be harnessed for social good.
Table 6 lists some recommendations of our results on how bots can be leveraged on for social good , based on possible avenues for interacting with the automated accounts. First, given that bots use more retweets and mentions than humans, and have high tweets per hour, bots can be used for menial tasks like announcements and distribution of information. Second, since bots have a star interaction network, they can be used for big announcements like disaster and crisis management without message distortion. A star network sends messages directly through interactions, hence the original message is preserved. However, the human’s hierarchal interaction network will distort the message as it passes through the tiers. Third, bots typically post content that matches their identity, they can be used to provide educational material about topics that people associate with certain profession. For example, a weather news bot can provide weather information. Lastly, since bots use more abusive and expletive terms than humans, instead of regulating toxic language itself, regulation can be focused on disallowing bots to use such toxic language, which would therefore reduce the amount of hyperbole and offense online. This regulatory recommendation draws on our observational study, and we suggests subsequent randomized controlled tests for validation before policy implementations.
@@
Further Investigations
In this paper, we primarily used an extensive dataset of Twitter data with tweets collected over four years. However, we have done other investigations to go beyond this dataset. In particular, we looked at bots across social media platforms and bot evolution.
Bots across Platforms
Are bots the same across different social media platforms? To answer that question, we perform a preliminary analysis towards investigating the similarities and differences of bots on different social media platforms.
We compared the psycholinguistic behavior of bots between Twitter and Telegram platforms with respect to the Coronavirus 2020-2021 pandemic. The Telegram dataset contained $\sim$7 million messages written by $\sim$335,000 users, and was segregated into three user groups: bots, disinformation spreaders and humans[100]. We extracted psycholinguistic cues using the NetMapper software for the Telegram messages in the same fashion as we did for the Twitter messages (see Methods: Comparison by Psycholinguistic Cues for details).
Figure 7 presents the differences between bots and humans for a subset of the psycholinguistic cues. We only compare the semantic and emotion cues because the structure of the Telegram platform is different from the Twitter platform, so the metadata cues do not map directly. This comparison shows that the pattern of cue usage between bots are humans are similar across both platforms, yielding generalizability to our overall study. In the same light, past work had shown that there was insignificant difference in the user metadata and post structure between tweets from Twitter and messages from Telegram[100]. Bots are generally similar in the usage of linguistic cues across social media platforms, which provides wider generalizability of our results.
Figure 7: Differences in the use of psycholinguistic cues between bots and humans for Twitter dataset and Telegram dataset. Both datasets are drawn from the Coronavirus2020-2021 event. Green cells show that humans use a larger number of the cue. Red cells show that bots use a larger number of the cue. * within the cells indicates there is a significant difference in the usage of the cue between bots and humans at the $p<0.05$ level.
Bot Evolution
Bots can evolve over time as adversarial actors realize that they are being detected and change their techniques [19]. Between 2021 and 2023, bots have enhanced their detection evasion techniques by adding four-character letter strings and Chinese proverbs to attempt to fool bot detection tools [101]. However, these techniques are not entirely novel. They are recombinations of social cybersecurity techniques like distract and distort techniques that were discovered in earlier studies [40, 32]. These techniques revolve around permuting the actions that platforms provide for content and interactions.
Bots also continue to evolve as new technologies, such as Generative AI technologies, come out. We used three open-sourced Large Language Models (LLMs) to generate tweets related to the coronavirus event. These models are: Meta Llama 3.1 8B Instruct, Phi-4, and Qwen 2.5 7B Instruct 1M. Each model generated 20 tweets. The Supplementary Material details the methodology and the prompts used for this experiment. Ideally, the use of generative AI technology would make the bot agents undetectable by current bot detection technology. That is, the use of LLMs would reduce the probability that the BotHunter algorithm classifies these tweets as originating from bot users. Across our experiments, the average BotHunter score retrieved was 0.69$\pm$0.15. This score is borderline on the 0.70 bot classification threshold that we used in this study, and is above the 0.50 threshold that many studies use [102, 34, 103, 104]. The huge standard deviation indicates that LLMs generate personas that are inconsistent with a single user type: sometimes LLMs generated tweets that look like humans, sometimes they do not.
Despite the evolution of bots, bot detectors are still able to identify users as bots. The same BotHunter tool used in this study had been deployed on data from 2023 to 2025 [16, 105, 106]. The tool classified 20% of the users as bot users, a percentage similar to that in our study. This shows that these bot detectors identify broad and comprehensive features. Moreover, these traditional bot types that do not employ generative AI still continue to exist, perhaps because the bot operator does not require generative technologies for his task. For example, a news aggregator bot that retweets a set series of news accounts can be programmed with heuristics. However, conversational bots will benefit from generative AI technologies and be able to carry out a more seamless conversation with the online community. Future work should examine the varied purposes of bots and whether evolution is required or whether evolution has occurred.
Future Work
Future work involves more detailed investigation into the nuances of bot behavior across platforms. It is also worth exploring how the content layout and the relationship structure of different platforms affect the users’ interaction dynamics. Future work can also take advantage of the temporal to analyze how bots evolve in behavior over time, and during crucial points in an event. The geographical dimension of our dataset further provides opportunity for a social geographical analysis on how bots differ across countries and regions.
@@
Bot evolution and bot detection are thus a “never-ending clash" [109], and sometimes bot accounts evolve faster than current known bot detection algorithms [70], presenting several opportunities in continual improvement of bot detection algorithms, specifically to be adaptable, faster, and more efficient. The increasing trends of using Large Language Models and Large Vision Models to create generated texts and deepfakes lend bots a helping hand in the construction of more believable narratives. These same generative technology are also used to construct offensive bots for humor [110]. However, current trends reflect that the use of such technologies is not very prevalent, for example, [75] only found one set of such botnet in their study, reflecting that bots are still relying on traditional techniques, likely because such heuristic-based techniques are easier and faster to deploy en masse.
Beyond simply detecting bot users in social media discourse, the next research agenda is to perform content-interaction impact analysis. Our work revealed one common interaction patterns bots use in their content communication. Hereupon arises an opportunity to characterize archetypes of interaction patterns, the best type of content to disseminate with each archetype, and the impact of the dissemination strategy on public opinion.
@@
After identifying which users are likely to be bots, one must differentiate the goodness of the bot and its function. This evaluation can be inferred from the bot’s content postings and relationship interactions. However, bots do not fall squarely in a spectrum of goodness; the lines of good and bad bots are blurred. In fact, bots can move between neutral in which they post messages that are not harmful, to bad, where they post politically charged and extremist messages [40, 111]. Herein lies an opportunity to construct a rubric to determine the goodness of the bot; this, though, is a complex task, for there are ethical and societal issues to consider. Bots can change their goodness, too. They may be supporting a certain cause initially, then making a swing to a different stance soon enough. This swing of support was witnessed during the coronavirus pandemic era, where bots change their stances towards the vaccination campaign, disseminating narratives of different stances during different periods [85, 8]. This opinion dynamics is important to the health of social discourse, and especially so when the bots require little conviction to change allegiances [8]. Another challenge involves identifying the type of bot, which can provide insight towards possible impact of the bot. For example, an Amplifier Bot that intensifies political messages could be intended to sow discord [24, 63].